OpenStudy (anonymous):

Use Fixed Point iteration to find the root. f(x)=x^4-x+0.2 near 1 , x0=1

OpenStudy (anonymous):

well we want to solve \[x^4-x+0.2=0\]we must isolate one the \(x\)'s in that expression

OpenStudy (anonymous):

\[x^4-x+0.2=0\]\[x=x^4+0.2\]or\[x=\sqrt[4]{x-0.2}\]which one will converge? what do u think?

OpenStudy (anonymous):

i should take derivative to check ??

OpenStudy (anonymous):

emm... just plug some values and see what happens u have 2 equations 1 \[x_{n+1}=x_n^4+0.2 \ \ \ \ \ x_0=1\] 2 \[x_{n+1}=\sqrt[4]{x_n-0.2} \ \ \ \ x_0=1\]

OpenStudy (anonymous):

second one will converge i think.

OpenStudy (anonymous):

Yup...i think second one is better. lets check it

OpenStudy (anonymous):

n=0\[x_{1}=\sqrt[4]{x_0-0.2}=\sqrt[4]{1-0.2}=0.9457 \]n=1\[x_{2}=\sqrt[4]{x_1-0.2}=\sqrt[4]{0.9457-0.2}=0.9292 \]n=2\[x_{3}=\sqrt[4]{x_2-0.2}=\sqrt[4]{0.9292-0.2}=0.924 \] oh converged very soon

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

lol...u playin football there?

OpenStudy (anonymous):

yup :P

OpenStudy (anonymous):

messi or ronaldo?

OpenStudy (anonymous):

...messi is the best ....haha...

OpenStudy (anonymous):

yeah he is :). can you help me with another problem ?

OpenStudy (anonymous):

np :)